OpenStudy (daniellelovee):

1. On a summer day, you take a road trip through Death Valley, California, in an antique car. You start out at a temperature of 21°C, but the temperature in Death Valley will reach a peak of 51°C. The tires on your car hold 15.6 L of nitrogen gas at a starting pressure of 249 kPa. The tires will burst when the internal pressure (Pb) reaches 269 kPa. Answer the following questions and show your work.

OpenStudy (daniellelovee):

• How many moles of nitrogen gas are in each tire? • What will the tire pressure be at peak temperature in Death Valley? • Will the tires burst in Death Valley? Explain. • If you must let nitrogen

OpenStudy (ichliebedich):

"How many moles of nitrogen gas are in each tire?" use the ideal gas law PV=nRT "What will the tire pressure be at peak temperature in Death Valley?" use P1V1/T1=P2V2/T2

OpenStudy (daniellelovee):

lol you totally copied that from another question like mine :p I see what you did there xD

OpenStudy (ichliebedich):

THE QUESTION IS TRICKY SO I search it

syed98 (syedmohammed98):

i think for the second one ichlie is right...

OpenStudy (daniellelovee):

what about the first one though?

OpenStudy (ichliebedich):

P = pressure (249 kPa = 2.4574 atm) V = 15.6 L n = moles T = Temp in Kelvin (21 deg C = 273.15 + 21 = 294.15 Kelvin) R = gas constant = 0.0821 L atm mol^-1 K^-1 i hope no one has this on other answers

OpenStudy (ichliebedich):

:/

syed98 (syedmohammed98):

well there are ways to find moles ..u can use number of moles = mass/molar mass but there's no mass given :(......u can also find it by using the formula V = moles * 24dm^3 but it can't be used on this one...so i m also confused....:l

OpenStudy (ichliebedich):

n = 1.5874 moles N2 n = 1.59 mol

OpenStudy (ichliebedich):

the answers are on the website you gave me

OpenStudy (somy):

what exactly do u not get here?

OpenStudy (daniellelovee):

nvm lol you were all correct I got it -.-'

OpenStudy (somy):

On a summer day, you take a road trip through Death Valley, California, in an antique car. You start out at a temperature of 21°C, but the temperature in Death Valley will reach a peak of 51°C. The tires on your car hold 15.6 L of nitrogen gas at a starting pressure of 249 kPa. The tires will burst when the internal pressure (Pb) reaches 269 kPa. Answer the following questions and show your work. lets see whats known T1= 21°C T2= 51°C V1= 15.6 L P1= 249 kPa Bursting pressure= 269 kPa so these are known to you now qs How many moles of nitrogen gas are in each tire? - We know that it can hold 15.6 L so we are going to use PV=nRT here, unknown is 'n' so lets make it as a subject of this formula n= PV/RT P1, V1, R (according to units of the rest of the members of formula) and T1 are known to you so put them into formula and find mole of nitrogen ********************************************** What will the tire pressure be at peak temperature in Death Valley? \[\frac{ P_1 }{ T_1 } = \frac{ P_2 }{ T_2 }\] you have P1, T1 and T2 given to you do find P2 using this formula ********************************************** Will the tires burst in Death Valley? Explain. The Pressure at which tire will burst is given to you compare it with P2 that you will find in 2nd section and see, if its less than bursting pressure- it will not burst, if its equal or more than that- it will burst ************************************************* Finish till here, then will move to this last one If you must let nitrogen gas out of the tire before you go, to what pressure must you reduce the tires before you start your trip? (Assume no significant change in tire volume.)

OpenStudy (somy):

owh lol didnt see ur response :3

OpenStudy (daniellelovee):

lol is ok thank you though ^.^

OpenStudy (somy):

no problem :)

OpenStudy (daniellelovee):

(can you guys share medals because you know I cant give more than 1 medal :))

OpenStudy (somy):

sure :D

OpenStudy (ichliebedich):

thanks

syed98 (syedmohammed98):

Thanks..